# Normal liquid 3He studied by Path Integral Monte Carlo with a parametrized partition function

### v1

@@ Abstract
We compute the energy per particle of normal liquid 3He in the temperature range $0.15-2$ K using Path Integral Monte Carlo simulations, leveraging a recently proposed method to overcome the sign problem – a long-standing challenge in many-body fermionic simulations. This approach is based on introducing a parameter $\xi$ into the partition function, which allows a generalization from bosons ($\xi=1$) to fermions ($\xi=-1$). By simulating systems with $\xi\geq 0$, where the sign problem is absent, one can then extrapolate to the fermionic case at $\xi=-1$. Guided by an independent particle model that uncovers non-analytic behavior due to the superfluid transition, which is moderated by finite-size effects, we develop a tailored extrapolation strategy for liquid 3He that departs from the extrapolation schemes shown to be accurate in those cases were quantum degeneracy effects are weak, and enables accurate results in the presence of Bose–Einstein Condensation and superfluidity for $\xi>0$. Our approach extends the previously proposed framework and yields energy per particle values in excellent agreement with experimental data.
@@ I Introduction
In the past few years, a novel method to overcome the sign problem for many-body fermionic simulations has been introduced. This approach is based on the consideration that the quantum statistical partition function of a system of bosons and fermions differs only by the value a single parameter, $\xi$, which assumes the value $\xi=1$ for bosons and $\xi=-1$ for fermions. [10, 11] Since the partition function turns out to be a continuous function of $\xi$, one can envision an extrapolation procedure to the fermion case, based on results in the $\xi\geq 0$ regime, where PIMC calculations are not affected by the sign problem and can be carried out with acceptable computational resources. This approach has been successfully applied to the study of the electron gas [10] and solvable fermion models. [12] In practice, one computes the average energy per particle $E$ (or any other observable) along an isotherm of temperature $T$ for several values of $\xi\geq 0$. These isotherms are used to compute how $\xi$ depends on $T$ for fixed energy $E$. These studies showed that simple fitting functions, usually quadratic polynomials, for $\xi_{E}(T)$ extrapolate very well to $\xi=-1$, thus providing the temperature $T$ at which the fermionic system attains energy $E$.
In this work, we investigate how this approach can be used to compute $E(T)$ in the case of normal liquid 3He. We find that the presence of Bose–Einstein Condensation (BEC) and the accompanying superfluid transition in the $\xi>0$ region results in a non-analytic behavior for the function $\xi_{E}(T)$ which, despite being smoothed out by finite-size effects, requires more care in extrapolating to $\xi=-1$ than in the case of previous studies. [12] After recalling the general theoretical framework, we will discuss an independent particle model that highlights the effect of BEC on the values of $E(T,\xi)$ along an isotherm and provides important clues on the general form of $\xi_{E}(T)$ in this case. Secondly, we will study finite-size effects in the same model, and investigate how they affect the extrapolation procedure. Finally, based on extensive PIMC simulations of $E(T,\xi)$, we apply our extrapolation method to compute $E(T)$ for 3He, finding a very good agreement with experimental data.
@@ II Theoretical framework
where $\beta=(k_{\mathrm{B}}T)^{-1}$, $k_{\mathrm{B}}$ is the Boltzmann constant, the sign ’$+$’ is for bosons and ’--’ for fermions, and $\mathbf{R}=\left(\mathbf{r}_{1},\mathbf{r}_{2},\dots,\mathbf{r}_{N}\right)$ is a compact notation for all the particle coordinates. In Eq. (1), $\rho(\mathbf{R},\mathbf{R}^{\prime},\beta)=\langle\mathbf{R}|e^{-\beta H}|\mathbf{R}^{\prime}\rangle$ is the canonical density matrix corresponding to the Hamiltonian $H$, and $\mathcal{P}\mathbf{R}=(\mathbf{r}_{\mathcal{P}(1)},\mathbf{r}_{\mathcal{P}(2)},\dots,\mathbf{r}_{\mathcal{P}(N)})$ denotes the position vectors with permuted labels.
@@
where $N_{\mathcal{P}}$ denotes the minimum number of times for which pairs of indices have to be interchanged in the current permutation $\mathcal{P}$ to recover the original (diagonal) order. [11, 12, 10, 13, 14]
In Eq. (2), the parameter $\xi$ interpolates continuously from bosons ($\xi$ = 1), to distinguishable particles ($\xi=0$), to fermions ($\xi=-1$), and can take any real value in ($\infty$- ∞,$\infty$). The energy of the system described by Eq. (2) can be obtained from standard statistical mechanics considerations as
@@
where the subscript means the $\xi$ variable is kept fixed. In order to estimate the energy in the case of fermionic systems, one can then exploit the fact that both $Z_{\xi}(T)$ and $E_{\xi}(T)$ are continuous functions of $T$ and $\xi$. By knowing their behavior in the $\xi\geq 0$ sector, where calculations can be performed without encountering the sign problem, one can extrapolate the value of the energy to $\xi=-1$. In fact, for $\xi\geq 0$ one can perform standard PIMC simulations [15] just by taking into account the different positive weights due to the factor $\xi^{N_{\mathcal{P}}}$ in Eq. (2).
There are two necessary conditions that can be used to guide the extrapolation to $\xi=-1$[11]
@@ III Independent-particle model
has been studied in Ref. 16, where it is shown that in an independent-particle model having Hamiltonian $H=\sum_{\mathbf{q}}\epsilon(\mathbf{q})\hat{N}_{\mathbf{q}}$ (where $\hat{N}_{\mathbf{q}}$ is the usual number operator), occupation numbers have the form
@@
where $\Omega$ is the volume of the system and the dispersion relation depends only on $q=|\mathbf{q}|$ due to spherical symmetry. We will assume in the following that $\epsilon(q>0)>0$ and that $\epsilon(q\to 0)=0$.
The symbol $N_{\mathrm{c}}$ in Eq. (9) represents the non-zero macroscopic number of particles in the ground state (i.e., the state with $\epsilon=0$) that signals the onset of BEC, which is characterized by the fact that $N_{\mathrm{c}}/N$ remains finite in the thermodynamic limit. [17] The presence of BEC in this system is related to the fact that for any $\xi>0$ the chemical potential has an upper bound, lest the occupation numbers of Eq. (8) become negative. Inspection of Eq. (8) shows that the chemical potential must be smaller than the critical value $\mu_{\mathrm{c}}=-T\log\xi$. Conversely, if one considers Eq. (9) as a function of $\xi$ along an isotherm, there is a critical value of the parameter $\xi$ after which $N_{\mathrm{c}}>0$ given by
$\xi_{c}$ $=\frac{\Omega}{N(2\pi)^{3}}\int_{0}^{\infty}\frac{4\pi q^{2}}{\text{exp}\left(\frac{\epsilon(q)}{T}\right)-1}\mathrm{d}q.$ (11)
@@
where $m=3.016$ a.m.u. is the mass of 3He. We notice that by tuning the two parameters, and $\alpha_{1}$ in particular, one can also obtain a free dispersion ($\alpha_{1}$=0, short dashed line in Fig. 1) or a Bogoliubov-like dispersion ($\sqrt{\alpha_{1}\alpha_{2}/\hbar^{2}}\sim 160$ [m/s]). Here the two parameters $\alpha_{1}$ and $\alpha_{2}$ are taken such that $\alpha_{2}=2.45$ and $\sqrt{\alpha_{1}\alpha_{2}/\hbar^{2}}\sim 302$ [m/s]. The studied dispersion is shown in Fig. 1 by a continuous line.
@@
We solved numerically Eqs. (9) and (10) fixing the density at $mN/\Omega=0.081$ g/cm3, and report our results in Fig. 2. Notice the presence of a cusp in the curves $E_{T}(\xi)$ in correspondence of the critical value $\xi_{\mathrm{c}}$. This behavior is due to the fact that the system undergoes BEC and hence a progressively larger macroscopic number of particles occupies the lowest-energy state, $\epsilon=0$, contributing to an abrupt reduction of the average energy and a non-analytic behavior of $E_{T}(\xi)$.
Figure 2: (a) Energies as a function of the $\xi$ parameter for the dispersion in Eq. (13) from Eqs. (9) and (10) for temperatures between $1$ K and $10$ K. Red stars indicate the energies obtained at $\xi_{\mathrm{c}}$ in Eq. (11). (b) Behaviour of $\xi_{E}(T)$ for a few selected energies. The latter are chosen at the critical points for $T=4,6,8$, and $10$ K. Notice that the slope change of $\xi_{E}(T)$ is located at these temperatures. (c) Reconstructed energies as function of $T$. Black dashed line: numerical solution. Blue crosses: extrapolation of Eq. (14) fitted using only data for $\xi\geq 0$. Orange circles: extrapolation of Eq. (6) fitted using only data for $\xi\geq 0$.
@@
We report in Fig. 2(b) the $\xi_{E}(T)$ functions, computed numerically from Eqs. (9) and (10), at energies corresponding to the critical points of the $E_{T}(\xi)$ curves (see Fig. 2(a)) for $T=4,6,8$, and $10$ K. One can see that the functions $\xi_{E}(T)$ are characterized by two regimes: there is a steep increase at low temperatures, until the critical point $\xi_{\mathrm{c}}$ is reached. Subsequently, the increase is less pronounced. This double regime reflects the fact that the system goes through the superfluid transition. As shown in Fig. 2(b), we found that the following function
$\begin{split}\xi_{E}(T)=&\left(a_{0}(E)+a_{2}(E)T^{2}+a_{3}(E)T^{3}\right)\Theta(T-T_{c})+\\ &\left(b_{0}(E)+b_{1}(E)T+b_{2}(E)T^{2}\right)\bar{\Theta}(T-T_{c}),\end{split}$ (14)
@@
provides a good fit for the $\xi_{E}(T)$ curves, where $T_{\mathrm{c}}$ is the temperature of the inflection point, $\Theta(T)$ is the Heaviside function and $\bar{\Theta}(T-T_{\mathrm{c}})=(1-\Theta(T-T_{\mathrm{c}}))$. Notice that the function of Eq. (14) has the form of Eq. (6) for small $T$, becoming a quadratic polynomial for $T>T_{\mathrm{c}}$. By solving $\xi_{E}(T)=-1$ one obtains the temperature $T$ where the fermion system attains energy $E$.
@@
In practice, one can compute $\xi_{E}(T)$ only for positive $\xi$, hence we re-fitted the parameters of Eq. (14) using only results in this region. We show in Fig. 2(c) (crosses), the results of the extrapolation to $\xi=-1$, compared to the actual value obtained by solving Eqs. (9) and (10) (dashed line). The agreement of the extrapolated energy and the actual energy is very good. In the same figure, we also show the result of the extrapolation to $\xi=-1$ using the a simple cubic form from Eq. (6), [12] fitted from the data at $\xi\geq 0$. In this case, the energy of the fermionic system is not accurately captured, and in some instances, the cubic polynomial fails to intersect the line $\xi=-1$.
IV Finite-size effects
In the case of PIMC simulations, the results of energy calculations are generally influenced by finite-size effects.

### v2

@@ Abstract
We compute the energy per particle of normal liquid 3He in the temperature range $0.15-2$ K using Path Integral Monte Carlo simulations, leveraging a recently proposed method to overcome the sign problem – a long-standing challenge in many-body fermionic simulations. This approach is based on introducing a parameter $\xi$ into the partition function, which allows a generalization from bosons ($\xi=1$) to fermions ($\xi=-1$). By simulating systems with $\xi\geq 0$, where the sign problem is absent, one can then extrapolate to the fermionic case at $\xi=-1$. Guided by an independent particle model that uncovers non-analytic behavior due to the superfluid transition, which is moderated by finite-size effects, we develop a tailored extrapolation strategy for liquid 3He that departs from the extrapolation schemes shown to be accurate in those cases were quantum degeneracy effects are weak, and enables accurate results in the presence of Bose–Einstein Condensation and superfluidity for $\xi>0$. Our approach extends the previously proposed framework and yields energy per particle values in good agreement with experimental data.
@@ I Introduction
In the past few years, a novel method to overcome the sign problem for many-body fermionic simulations has been introduced. This approach is based on the consideration that the quantum statistical partition function of a system of bosons and fermions differs only by the value a single parameter, $\xi$, which assumes the value $\xi=1$ for bosons and $\xi=-1$ for fermions. [10, 11] Since the partition function turns out to be a continuous function of $\xi$, one can envision an extrapolation procedure to the fermion case, based on results in the $\xi\geq 0$ regime, where PIMC calculations are not affected by the sign problem and can be carried out with acceptable computational resources. This approach has been successfully applied to the study of the electron gas [10] and solvable fermion models. [12] In practice, one can directly extrapolate the energy (or other observables such as the pair correlation functions) obtained at different $\xi\geq 0$. But in a more general framework, [12] that will also be used in this work, one can first compute the average energy per particle $E$ (or, in principle, also other observables) along an isotherm of temperature $T$ for several values of $\xi\geq 0$. These isotherms are then used to compute how $\xi$ depends on $T$ for fixed energy $E$. It has been shown that simple fitting functions, usually quadratic polynomials, for $\xi_{E}(T)$ extrapolate very well to $\xi=-1$, thus providing the temperature $T$ at which the fermionic system attains energy $E$[12] However, the application of this extrapolation method so far has been performed under condition of weak quantum degeneracy effect. [10]
In this work, we investigate how this approach can be used to compute $E(T)$ in a case of normal liquid 3He, where quantum degeneracy is significant. We find that the presence of Bose–Einstein Condensation (BEC) and the accompanying superfluid transition in the $\xi>0$ region results in a non-analytic behavior for the function $\xi_{E}(T)$ which, despite being smoothed out by finite-size effects, requires more care in extrapolating to $\xi=-1$ than in the case of previous studies. [12] After recalling the general theoretical framework, we will discuss an independent particle model that highlights the effect of BEC on the values of $E(T,\xi)$ along an isotherm and provides important clues on the general form of $\xi_{E}(T)$, that hold for both the fully spin-polarized and spin-unpolarized cases. Secondly, we will study finite-size effects in the same model, and investigate how they affect the extrapolation procedure. Finally, based on extensive PIMC simulations of $E(T,\xi)$, we apply our extrapolation method to compute $E(T)$ for 3He, finding a good agreement with experimental data and previous theoretical results.
@@ II Theoretical framework
where $\beta=(k_{\mathrm{B}}T)^{-1}$, $k_{\mathrm{B}}$ is the Boltzmann constant, the sign ’$+$’ is for bosons and ’--’ for fermions, and $\mathbf{X}=\left(\mathbf{r}_{1},s_{1};\mathbf{r}_{2},s_{2};\dots,\mathbf{r}_{N},s_{N}\right)$ is a compact notation for all the particle coordinates, that is positions $\mathbf{r}_{i}$ and spins $s_{i}$. In Eq. (1), $\rho(\mathbf{X},\mathbf{X}^{\prime},\beta)=\langle\mathbf{X}|e^{-\beta H}|\mathbf{X}^{\prime}\rangle$ is the canonical density matrix corresponding to the Hamiltonian $H$, and $\mathcal{P}\mathbf{X}=(\mathbf{r}_{\mathcal{P}(1)},\sigma_{\mathcal{P}(1)};\mathbf{r}_{\mathcal{P}(2)},\sigma_{\mathcal{P}(2)};\dots,\mathbf{r}_{\mathcal{P}(N)},\sigma_{\mathcal{P}(N)})$ denotes the position vectors with permuted labels.
@@
where $N_{\mathcal{P}}$ denotes the minimum number of times for which pairs of indices have to be interchanged in the current permutation $\mathcal{P}$ to recover the original (diagonal) order. [11, 12, 10, 13, 14] In the following, we will be concerned with a spin-independent Hamiltonian $H$, hence it is convenient to write the states in the factorized form $|\mathbf{X}\rangle=|\mathbf{R}\rangle|\mathbf{S}\rangle$, where $\mathbf{R}$ is a shorthand for all the positions $\mathbf{r}_{i}$ and $S$ denote all the single-particle spins; the density matrix can then be written
$\rho(\mathbf{X},{\cal P}\mathbf{X},\beta)=\rho(\mathbf{R},{\cal P}\mathbf{R},\beta)\langle S|{\cal P}S\rangle.$ (3)
Equation (3) shows that the density matrix $\rho$ is zero, unless the permutation operator ${\cal P}$ is such that it does not exchange particles with different spins. Considering particles with spin $\frac{1}{2}$, this requirement implies ${\cal P}={\cal P}_{-}{\cal P}_{+}$, where ${\cal P}_{-}$ is a permutation among particles with negative spin, and ${\cal P}_{+}$ is a permutation among particles with positive spin. This is equivalent to saying that since the single-particle spin is a good quantum number, particles of different spins are distinguishable [15] and the total spin $S$ is constant. This property enables simulations to be performed at different polarizations, by randomly assigning a fraction $\phi$ of the particles to positive spin, and – correspondingly – a fraction $1-\phi$ of the particles to negative spin, resulting in polarization $P\equiv S/N=2\phi-1$ During the Monte Carlo calculation, one independently samples the permutation among particles of positive or negative spin.
In Eq. (2), the parameter $\xi$ continuously interpolates from bosons ($\xi$ = 1), to distinguishable particles ($\xi=0$), to fermions ($\xi=-1$), and can take any real value in ($\infty$- ∞,$\infty$). The energy of the system described by Eq. (2) can be obtained from standard statistical mechanics considerations as
@@
where the subscript means the $\xi$ variable is kept fixed. In order to estimate the energy in the case of fermionic systems, one can then exploit the fact that both $Z_{\xi}(T)$ and $E_{\xi}(T)$ are continuous functions of $T$ and $\xi$. By knowing their behavior in the $\xi\geq 0$ sector, where calculations can be performed without encountering the sign problem, one can extrapolate the value of the energy to $\xi=-1$. In fact, for $\xi\geq 0$ one can perform standard PIMC simulations [16] just by taking into account the different positive weights due to the factor $\xi^{N_{\mathcal{P}}}$ in Eq. (2). Notice that in the ”bosonic” ($\xi>0$) regime, one is in general simulating a system that is a mixture of different bosons, one for each of the $12s+$ 1 values of the spin of the limiting fermionic particles. Only in the case of a fully polarized system ($P=1$) does the bosonic sector correspond to actual pure bosons.
In the case of a spin-independent Hamiltonian, there are two necessary conditions that can be used to guide the extrapolation to $\xi=-1$[11, 17]
@@
While the extrapolation approach of Eq. (5) has already been used to study observables different from the energy per particle (e.g., the pair correlation function, [11]), we stress that the possibility of extending the present approach based on the extrapolation of Eq. (7[12] to study other observables is feasible and will be investigated in future studies.
@@ III Independent-particle model
has been studied in Ref. 18, where it is shown that in an independent-particle model having Hamiltonian $H=\sum_{\mathbf{q}}\epsilon(\mathbf{q})\hat{N}_{\mathbf{q}}$ (where $\hat{N}_{\mathbf{q}}$ is the usual number operator), occupation numbers have the form
@@
where $\Omega$ is the volume of the system and the dispersion relation depends only on $q=|\mathbf{q}|$ due to spherical symmetry. We will assume in the following that $\epsilon(q>0)>0$ and that $\epsilon(q\to 0)=0$. The symbol $N_{\mathrm{c}}$ in Eq. (10) represents the non-zero macroscopic number of particles in the ground state (i.e., the state with $\epsilon=0$) that signals the onset of BEC, which is characterized by the fact that $N_{\mathrm{c}}/N$ remains finite in the thermodynamic limit. [19] The presence of BEC in this system is related to the fact that for any $\xi>0$ the chemical potential has an upper bound, lest the occupation numbers of Eq. (9) become negative. Inspection of Eq. (9) shows that the chemical potential must be smaller than the critical value $\mu_{\mathrm{c}}=-T\log\xi$. Conversely, if one considers Eq. (10) as a function of $\xi$ along an isotherm, there is a critical value of the parameter $\xi$ after which $N_{\mathrm{c}}>0$ given by
$\xi_{\mathrm{c}}$ $=\frac{\Omega}{N(2\pi)^{3}}\int_{0}^{\infty}\frac{4\pi q^{2}}{\text{exp}\left(\frac{\epsilon(q)}{T}\right)-1}\mathrm{d}q.$ (12)
@@
Figure 2: Independent–particle model. (a) Energies as a function of the $\xi$ parameter for the dispersion in Eq. (16) and for $P=1$ (fully polarized case) from Eqs. (10) and (11) for temperatures between $1$ K and $10$ K. Red stars indicate the energies obtained at $\xi_{\mathrm{c}}$ in Eq. (12). (b) Behaviour of $\xi_{E}(T)$ for a few selected energies for $P=1$. The latter are chosen at the critical points for $T=4,6,8$, and $10$ K. Notice that the slope change of $\xi_{E}(T)$ is located at these temperatures. (c) and (d): Same of (a) and (b) in the case of $P=0$ (fully unpolarized system).
Equations (10) and (11) have been presented, for the sake of simplicity, neglecting the spin variable. In the case of spin-independent Hamiltonians, the model can be straightforwardly generalized since, as discussed before, the $z$ component of the spin is a conserved quantum number. Therefore, in a single-particle approach we can consider our system as a mixture of particles of different spins; each spin component will be described by a pair of equations analogous to (10) and (11). Specializing to the case of 3He, we will then have a fixed number of particles with spin up $N_{\uparrow}$ and a fixed number of particles with spin down $N_{\downarrow}$ such that $N=N_{\uparrow}+N_{\downarrow}$, with corresponding chemical potentials $\mu_{\uparrow}$ and $\mu_{\downarrow}$, as well as energies $E_{\uparrow}$ and $E_{\downarrow}$. The polarization $P$ of the system will be then constant and equal to
$P=\frac{N_{\uparrow}-N_{\downarrow}}{N_{\uparrow}+N_{\downarrow}},$ (14)
and the total energy per particle, $E/N$, will be given by
$\frac{E}{N}=\frac{N_{\uparrow}}{N}E_{\uparrow}+\frac{N_{\downarrow}}{N}E_{\downarrow}.$ (15)
Notice that, according to Eq. (12), the various spin components will undergo BEC separately, that is the function $E_{T}(\xi)$ will display two non-analytic points, unless the system is unpolarized ($P=0$, hence $\xi_{\mathrm{c},\uparrow}=\xi_{\mathrm{c},\downarrow}$) or fully polarized ($P=-1$ or $P=1$).
@@
where $m=3.016$ a.m.u. is the mass of 3He. We notice that by tuning the two parameters, and $\alpha_{1}$ in particular, one can also obtain a free dispersion ($\alpha_{1}$=0, short dashed line in Fig. 1) or a Bogoliubov-like dispersion ($\sqrt{\alpha_{1}\alpha_{2}/\hbar^{2}}\sim 160$ [m/s]). Here the two parameters $\alpha_{1}$ and $\alpha_{2}$ are taken such that $\alpha_{2}=2.55$ and $\sqrt{\alpha_{1}\alpha_{2}/\hbar^{2}}=302$ [m/s]. The studied dispersion is shown in Fig. 1 by a continuous line. We solved numerically Eqs. (10) and (11) fixing the density at $mN/\Omega=0.081$ g/cm3, and report our results in Fig. 2 for the fully polarized (panels (a) and (b)) and unpolarized (panels (c) and (d)) systems. Notice the presence of a cusp in the curves $E_{T}(\xi)$ in correspondence of the critical value $\xi_{\mathrm{c}}$ for both the $P=1$ and $P=0$ cases. This behavior is due to the fact that the system undergoes BEC and hence a progressively larger macroscopic number of particles occupies the lowest-energy state, $\epsilon=0$, contributing to an abrupt reduction of the average energy and a non-analytic behavior of $E_{T}(\xi)$. In particular, we find that the value of $\xi$ at the critical point is twice as large in the $P=0$ regime than in the $P=1$ case (see Figs. 2(c) and (a)). Additionally, the behaviour of $E_{T}(\xi)$ is steeper in the polarized case than in the unpolarized case (see Fig. 8 in Appendix A). As a consequence, one has $E^{P=0}_{T}(\xi=1)>E^{P=1}_{T}(\xi=1)$ and, conversely, $E^{P=0}_{T}(\xi=-1)<E^{P=1}_{T}(\xi=-1)$. The energies of the polarized and unpolarized cases are the same at $\xi=0$, that is $E^{P=0}_{T}(\xi=0)\equiv E^{P=1}_{T}(\xi=0)$.
@@
We report in Figs. 2(b) and 2(d) the $\xi_{E}(T)$ functions, computed numerically from Eqs. (10) and (11), at energies corresponding to the critical points of the $E_{T}(\xi)$ curves (see Figs. 2(a) 2(c) respectively) for $T=4,6,8$, and $10$ K. One can see that these functions are characterized by two regimes: there is a steep increase at low temperatures, until the critical point $\xi_{\mathrm{c}}$ is reached. Subsequently, the increase is less pronounced. This double regime of $\xi_{E}(T)$ reflects the fact that the system goes through the superfluid transition. The effect of a non-zero polarization is to flatten the $\xi_{E}(T)$ curves. Figures 2(b) and 2(d), also report a fit the numerical data according to
$\begin{split}\xi_{E}(T)=&\left(a_{0}(E)+a_{2}(E)T^{2}+a_{3}(E)T^{3}\right)\Theta(T-T_{\mathrm{c}})+\\ &\left(b_{0}(E)+b_{1}(E)T+b_{2}(E)T^{2}\right)\bar{\Theta}(T-T_{\mathrm{c}}),\end{split}$ (17)
@@
where $T_{\mathrm{c}}$ is the temperature of the inflection point, $\Theta(T)$ is the Heaviside function and $\bar{\Theta}(T-T_{\mathrm{c}})=(1-\Theta(T-T_{\mathrm{c}}))$. As can be seen, these equations provide a good fit for the $\xi_{E}(T)$ curves. Notice that the fitting function of Eq. (17) has the form of Eq. (7) for small $T$, becoming a quadratic polynomial for $T<T_{\mathrm{c}}$. From the curves of Eq. (17) one obtains the temperature $T$ where the fermion system attains energy $E$ by solving $\xi_{E}(T)=-1$.
The fitting function in Eq. (17) captures the two-regimes observed in the independent-particle model, and emphasizes the role of the BEC transition temperature $T_{c}$, which always occurs at some $\xi>0$ in realistic systems independently on the specific form of the interaction. However, in the free-particle case, the BEC transition is hardly apparent in the $\xi_{E}(T)$ curves (see also appendix B) and a simple polynomial function is enough to properly describe its shape. [12]. In any case, Eq. (17) is robust enough to handle all such cases automatically. In theory, the polynomial expression in Eq. (7), with its infinite number of terms, is highly general and, in theory, capable of fitting any complicated function. Nevertheless, in actual calculations where only a limited number of points can be computed, Eq. (17) provides a more physically grounded expression to guide the extrapolation.
Figure 3: Independent-particle model. (a) Reconstructed energies as function of $T$ for $P=1$. Black dashed line: numerical solution. Blue crosses: extrapolation of Eq. (17) fitted using only data for $\xi\geq 0$. Orange circles: cubic extrapolation of Eq. (7) fitted using only data for $\xi\geq 0$. (b) Same of (a) for the unpolarized ($P=0$) case. We notice that the $P=0$ uncertainties are smaller than those corresponding to $P=1$, and that in both cases the magnitude of the error bar decreases by increasing the temperature.
@@
In practice, one can compute $\xi_{E}(T)$ only for positive $\xi$, hence we re-fitted the parameters of Eq. (17) using only results in this region. We show in Figs. 3(a) and 3(b) for the fully polarized and unpolarized cases respectively, the results of the extrapolation to $\xi=-1$ (blue crosses), compared to the actual value obtained by solving Eqs. (10) and (11) (black dashed line). The agreement between the extrapolated and actual energies is excellent.. However, at lower energies, the critical value of $\xi$ shifts closer to $\xi=0$, resulting in a larger uncertainty in the extrapolation. In the same figures, we also show the result of the extrapolation to $\xi=-1$ using a simple cubic form from Eq. (7), [12] fitted from the data at $\xi\geq 0$. In this case, the energy of the fermionic system is not accurately captured, and in some instances, the cubic polynomial fails to intersect the line $\xi=-1$.
Figure 4: Independent-particle model. (a) Comparison of the energy and its derivative with respect to $\xi$ for $T=6$ K for the bulk system (black points and black short-dashed line) and the finite size system (blue points and blue dashed-line) in the fully polarized case ($P=1$). The long-dashed line corresponds to the Fermi level. (b) Comparison of $\xi_{E}(T)$ for $E=E_{T=6}(\xi_{\mathrm{c}})$ K for the bulk system (black circles) versus the finite system (blue triangles). This value for the energy is chosen at the $\xi_{\mathrm{c}}$ of the bulk system in (a). (c) and (d) Same as (a) and (b) respectively but for $P=0$.
III.1 Finite-size effects